A force of 577 N acts for 3.99 m over 4.55 s. That force is the product of 708 J of energy to the system that is supplied for 1.20 s. What is the efficiency of the system?


1
Expert's answer
2021-04-28T09:54:59-0400

Explanations & Calculations


  • The efficiency is defined as

"\qquad\qquad
\begin{aligned}
\small Eff&=\small \frac{Work \,\,output}{Work\,\, input}\times 100

\end{aligned}"

  • Dividing both numerator and denominator yields another form for efficiency with power.

"\qquad\qquad
\begin{aligned}
\small Eff&=\small \frac{Work\,output/t}{Work \, input/t}\times100\\
&=\small \frac{Power \,output}{Power \,input}\times 100
\end{aligned}"

  • When power is given in each situation, this form can be used


  • Input power is

"\qquad\qquad
\begin{aligned}
\small &=\small \frac{708J}{1.20s}=590W
\end{aligned}"

  • Power output from the system is

"\qquad\qquad
\begin{aligned}
\small &=\small \frac{work}{time}=\frac{577N\times 3.99m}{4.55s}=505.98 W
\end{aligned}"

  • Then,

"\qquad\qquad
\begin{aligned}
\small Eff&=\small \frac{505.98\,W}{590\,W}\times 100\\
&=\small \bold{85.76\%}
\end{aligned}"